I need help with this problem 1/3(z+2)=1/9z
Answer by stanbon(75887)

1/3(z+2)=1/9z
----
Multiply both sides by 9 to get:
3(z+2) = z
---
Modify:
3z + 6 = z
2z = -6
z = -3
============
Cheers,
Stan H.
============